-4x + 5y = 11
y = 4x - 1

Answer by ikleyn(53765)

.

The second equation expresses  y  via  x.

Substitute this expression into the first equation.   You will get then


    -4x + 5*(4x-1) = 11.


Simplify and find x

    -4x + 20x - 5 = 11

    16x = 11 + 5

    16x = 16

      x = 1.


To find y, substitute the found value of x= 1 into the second equation

    y = 4*1 - 1 = 3.


ANSWER.  x= 1;  y= 3.

Solved.

It is how the substitution method works in this case.